GIF Lookbook

  • 01/07/2013 17:26: HVscan source ON (normal gas mixture) (dropbox)
  • 01/07/2013 18:24: HVscan source OFF (normal gas mixture) (dropbox)
  • 03/07/2013 15:06: HVscan source ON ARGON (dropbox)
  • 18/07/2013 15:47: changed HV cable 6.2 to 5.2 (Module ID: see sticker on the back of the module)
  • 18/07/2013 15:30: watertank was full, made it empty. Expected lower humidity
  • 19/07/2013 16:30: HVscan source OFF (normal gas mixture) (dropbox)
  • 19/07/2013 20:15: HVscan source ON (normal gas mixture) (dropbox)
  • 24/07/2013 17:20: Argon flow on (all gaps/ch 2l/h) T=21.5C, p=961mbar, humidity=51%, airco-set: 21C SOURCE ON
  • 26/07/2013 17:27: HVscan source ON ARGON (dropbox)
  • 26/07/2013 19:42: HVscan source OFF ARGON (dropbox)
  • 01/10/2013 new GIF software ready with averaging
  • 01/10/2013 21:28: HVscan source ON ARGON (dropbox)
  • 02/10/2013: very high humidity, water tank emptied
  • 26/11/2013: Setup stopped , whole trolley moved to 904
  • 30/11/2013: New RE4 gaps selected from various Batches see attached file : https://twiki.cern.ch/twiki/pub/Main/GIFTest/30112013_selection_gap_qc.xlsx
  • 28/01/2014: Dose measurements done.
  • 31/01/2014: New sets of chambers and gaps installed at GIF.
  • 06/02/2014 15:15: HVscan source ON ARGON (dropbox) T=20.01, RH=43% - Gas flow started 5/02 7:30 pm flow/channel ~3l/hour
  • 7/02/2014: Standard gas mixture Total Input Gas flow : 25l/hour - Flow / gas channel : ~ 3 l/hour
  • 10/02/2014 : 5:00 pm Additional humidier installed, set point =40% 5:30 pm HV scan (SOURCE ON, CMS mixture) started 9:30 pm HV scan stopped at 6 KV, connection between the mainframe and the detectors broken
  • 11/02/2014 11:30 pm SOURCE OFF 12:00 HV scan ( SOURCE OFF, CMS mixture) started 6:00 pm HV scan stopped Meteo station stopped updating the data file
  • 12/02/2014 11:57 SOURCE ON meteo station seems to work properly 12:56 HVscan (SOURCE ON, CMS mixture) Started , HV scan unsuccesfull , problems with meteo and main frame
  • 20/02/2014 23:18: HVscan source ON (normal gas mixture) (dropbox)
  • 27/02/2014 meteo station down, installed a new meteo, Ghent-030 tripping --> 3polar connector of the red extension cable gives sparks Ghent-030 disconnected
  • 27/02/2014 18:45: HVscan source ON (normal gas mixture, meteo 904) (dropbox)
  • 28/02/2014 14:40: HVscan source OFF (normal gas mixture, meteo 904) (dropbox)
  • 17/03/2014 problems with the computer, we removed the meteo station from the computer located on the right of the table ( pcta1gs27) and connected to the second computer located on the left of the table: PC Id : pcdt1gs31.cern.ch We installed on the new computer teamviewer , ID :111 699 114 with a fixed password : the same as in the previous computer. The data from the meteo are now written in the folder : C:\Program Files\Oregon Scientific\Weather OS\DATA Also we installed two additional ( T& RH ) sensors which are wireless connected with the main main station Now we have data from 3 sensors : The sensor on the main meteo unit located inside the gas room called sensor-0 one inside the RPCs tent : sensor-1 one close to the GRPCs : sensor-2 We connected the Ghent-030 chamber to the HV ( Red extension cable repaired)
  • 17/03/2014 19:31: HVscan source OFF (normal gas mixture, meteo 904) (dropbox)
  • 18/03/2014 16:49: HVscan source ON (normal gas mixture, meteo 904) (dropbox)
  • 21/03/2014 Meteo data are now written in C:\Documents and Settings\LocalAccount\Desktop\MeteoGIF in order to make the folder shareable. The sensor-1 appears inactive. Sensor-0 apparently produces measurements with a wrong timestamp.
  • 23/03/2014 METEO STATION: now sensor-0 is ok. Sensor-1 is still inactive. Actually sensor-2 is used to produce data for the HV correction.
  • 01/04/2014 17:30 source OFF, baterries of the sensor-1 changed, sensor-1 moved from the tent ouside, T, RH are fake, Main unit reset, Datalogger update configured to every 15 min. Both sensors are seen by the meteo, 18:15 : sensor-1 moved inside the tent, 18:20 Source ON
  • 09/04/2014 GHENT-030 removed from GIF, tested in 904 lab: HVscan OK (click here), currently under stabilitytest at 904
  • 09/04/2014: wireless sensor communication problems due to the radiation and concrete blocks. Lead blocks placed in front of the sensor in order to reduce the radiation, but the use of too much lead leads to bad signal strength. Tried to find the optimal position of the sensors, currently without success
  • 11/04/2014 Base unit meteo station moved inside the concrete blocks, behind the source at 2m height. Sensors behind lead blocks. Tested on 11-12-13 and 14 April: connection is stable, meteo station is working perfectly.
  • 15/04/2014: GHENT-030 for SM, new chamber for GIF will be installed.
  • 19/04/2014 : New chamber in GIF : RE4_3_CERN_004, Assembly Date: 30-NOV-12 Bottom Gap:KODEL-CMS-RE4-3-B003 Top Wide Gap :KODEL-CMS-RE4-3-TW020 Top Narrow Gap :KODEL-CMS-RE4-3-TN004 Distribution Board 15 FEB A :10037,FEB B:10041,FEB C: 10048 Adapter Board A1: 648,Adapter Board A2: 649, Adapter Board B1: 646,Adapter Board B2 647,Adapter Board C1 664, Adapter Board C2 665, Shipment Date to CERN: 18-APR-14, Reception Date at CERN: 18-APR-14
  • 30/04/2014 STABILTY STARTED with all gaps
  • 02/05/2014 Source off from 10:50 to 16:20
  • 06/05/2014 6:30 am scheduled Power cut , computer off, gas off, source off but main frame ON (HV~2kV on all the gaps ), stability not stopped
  • 08/05/2014 9:20 am stability program stopped because computer in 904 went off 10:20 pm stability program restarted HV~ 2KV on all the gaps (computer in 904 restarted)
  • 09/05/2014 9:30 am Severiano switch on the computer meteo started again but he didn't check the gas , gas was still off, 11:00 am stability set HV at 9.5 KV because meteo station sent again data but gas is still off
  • 16/05/2014 6:50 pm Nicolas went to GIF and realized that gas was not circulating he found the gas room without power, all the equipments in the gas room are off, computer controlling the mass flow meters is off, etc. , no any gas circulating gas kill Unit is also off. In the rest of the building there is power, Source is ON, Grpcs equipments in the source area are on. Air condition inside the RPC- tent was off. Nicolas switched it on Main frames in the control rooms were ON ,computer of Lyon in the control room is ON. Concerning the RPCs, no gas kill signal on the frame (because the electronic Unit of the gas kill if OFF), Nicolas found several channels tripped, he switched off the HV for the rest of the channels,HV stopped for all gaps at 7:00 pm STABILITY STOPPED at 07:10 pm , gas piquet called
  • 20/05/2014 Gas back The settings of the mass flow meters : Freon : C2H2F4 20 l/h, 20/21.05 = 95.0 % Iso C4H10 1 l/h 1/21.05 = 4.976 % SF6 0.05 l /h 0.05/21.05 = 0.024 % All channels are bubbling humidifier 4/6 40% humidity Ambient humidity in the tent 50%, T =19.7 degrees MAIN FRAME ON
  • 23/05/2014 11:03 am HV scan source OFF, standard mixture 10:38 pm HV scan source ON , standard mixture
  • 24/05/2014 STABILITY STARTED for all the gaps
  • 27/05/2014 10:40 am SOURCE OFF
  • 30/05/2014 5:30 pm found alarm in the control room , the source is again ON. Currents checked, HV =9500 except for the RE4- CERN-004-bot, HV=9350)
  • 07/06/2014 STABILITY STOPPED Because expiration of duration (14 days by default)
  • 11/06/2014 9h30 am , Martin found the source OFF and in alarm (buzzer). Investigation is currently ongoing, they will restart the source as soon as the problem is understood.
  • 11/06/2014 Source back ON at 10h44 am
  • 11/06/2014 01:00 pm STABILITY RESTARED ( expiration is 20 days)
  • 14/06/2014 7:40 am power cut, source went OFF, power in the gas room off --> no gas circulation, computer off, , air condition off, no meteo data, From log file of stability software connexion with main frame has been lost no any data until 5:20 pm, at 5:20 pm connexion with main frame established, HV= 9.5 KV meteo station still off line , the program remains stable at the last applied voltage because meteo file is not updated
  • 16/06/2014 12:30 pm , Nicolas set the HV OFF and STABILITY Stopped, gas piquet called, 02:00 pm gas ON , 02:45 pm source ON
  • 18/06/2014 3:00 pm Stability started , air condition off, new version of the software deployed including Temperature correction (T ref = 20 degrees) , meteo on, gas on,
  • 19/06/2014 2:45 pm power cut in the gas room simulated, Hv off for all the channels, tests with stability prgm , 3:00 pm air condition again ON, 6:00 pm fake gas kill The gas is circulating normally, the box is generating continuously the gas kill signal, Gas piquet called
  • 20/06/2014 Gas kill electronics solved, HVscan source ON on all gaps
  • 21/06/2014 Monitoring started (all gaps 9.5 kV), temperature correction included
  • 23/06/2014 Source off due to installation equipment at GIF?
  • 17/07/2014 P-T correction formula updated for HVscan and monitoring

-- GabriellaPugliese - 15 Dec 2013

Topic attachments
I Attachment History Action Size Date Who Comment
Unknown file formatxlsx 201404091813_hvscan_CMS-RE4-3-GHENT-030-TOP.xlsx r1 manage 13.8 K 2014-04-10 - 12:29 NicolasZAGANIDIS HVscan GHENT-030 at 904 lab
Edit | Attach | Watch | Print version | History: r23 < r22 < r21 < r20 < r19 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r23 - 2014-07-17 - JanEysermans
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback